OpenStudy (anonymous):

What is the value of x, when 10(x + 2) = 5(x + 8)?

jhonyy9 (jhonyy9):

like a first step 10*x +10*2 = 5*x +5*8

OpenStudy (unklerhaukus):

10(x + 2) = 5(x + 8) the first step, is to divide both sides of the equation by 5 2(x + 2) = x + 8 then, distribute on the left hand side 2x + 4 = x + 8 now, take away x+4 from both sides of the equation x =